OpenStudy (anonymous):

Given the function f(x)=(x-9)^2 where x is less than or equal to 9 , find f^-1(x) and its domain. The domain of is .

OpenStudy (turingtest):

y=(x-9)^2 \[\sqrt{y}=x-9\]\[x=\sqrt{y}+9\]so the inverse of f(x) is\[f^{-1}(x)=\sqrt{x}-9\] The domain is the set of all x for which f(x) is defined. We are told that \[x \le 9\] and if \[x<0\] the radical gives an imaginary number, so the range is\[0<x \le 9\] or in interval notation:\[(0,9]\]

OpenStudy (turingtest):

sorry, I meant the DOMAIN is \[0\le x \le 9\] or \[[0,9]\]

OpenStudy (turingtest):

x=0 is okay since it doesn't yield an undefined or imaginary number, so I used closed brackets.

OpenStudy (anonymous):

i tried this, but its all wrong according to my online thing :(

OpenStudy (turingtest):

you got the same answer as me?

OpenStudy (anonymous):

i got sqrt(x)+9 but i typed yours in and both were wrong as well as the interval notation

OpenStudy (turingtest):

yeah my bad, should be sqrt(x)+9 but that doesn't change the answer: [0,9] if your online thing says this interval is wrong I am suspicious of it. Still, don't give me a medal (so your post still says "needs a good answer") and see if anyone else comes up with something different. I can't see how my answer is wrong.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i have another question on the newfeed if u could please help thanks for ur time